# Causal3D: A Comprehensive Benchmark for Causal Learning from Visual Data

### v1

@@ Abstract
True intelligence hinges on the ability to uncover and leverage hidden causal relations. Despite significant progress in AI and computer vision (CV), there remains a lack of benchmarks for assessing models’ abilities to infer latent causality from complex visual data. In this paper, we introduce Causal3D, a novel and comprehensive benchmark that integrates structured data (tables) with corresponding visual representations (images) to evaluate causal reasoning. Designed within a systematic framework, Causal3D comprises 19 3D-scene datasets capturing diverse causal relations, views, and backgrounds, enabling evaluations across scenes of varying complexity. We assess multiple state-of-the-art methods, including classical causal discovery, causal representation learning, and large/vision-language models (LLMs/VLMs). Our experiments show that as causal structures grow more complex without prior knowledge, performance declines significantly, highlighting the challenges even advanced methods face in complex causal scenarios. Causal3D serves as a vital resource for advancing causal reasoning in CV and fostering trustworthy AI in critical domains.
Machine Learning, ICML
Figure 1: The proposed Causal3D dataset. We display 8 real-world scenes (11 hypothetical scenes are in the Appendix A). We focus on 3 scenes: springs, parabolas, and water flow. 1) The blue block represents multi-view images of each scene, offering four different views. The first row shows virtual backgrounds, while the second row shows real backgrounds, with the same view in each column; 2) The green block provides textual descriptions; 3) The yellow block represents the causal graphs for each scene, along with the meanings of each variable in the graphs; 4) The pink block shows the structural equations (i.e., functions describing causal relations) for each scene. The bottom row briefly presents an overview of the remaining 5 real-world scenes and their corresponding causal graphs, including reflection, seesaw, convex lens, magnet, and pendulum. Detailed information on these 5 scenes can be found in the Appendix A.
@@ 1 Introduction
Some existing datasets have been involved in causal studies in visual AI systems. However, these datasets often have significant limitations. For instance, oversimplified 2D hypothetical datasets (Yang et al., 2020) fail to capture the richness and complexity of real-world environments. Similarly, domain-specific datasets like the CelebA face dataset (Liu et al., 2015) are not designed for causal reasoning and lack the structural diversity required for comprehensive explorations. Recent datasets developed for vision- and multimodal-language models (VLMs/MLMs) (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Mao et al., 2022; Tung et al., 2024) have improved in complexity and realism but remain limited in explicitly representing causality and in offering diverse causal relations. The absence of clearly defined causal relations within visual representations, along with the lack of tabular records that are tightly aligned with these representations to provide guidance, makes such physics-aware VLMs/MLMs datasets suboptimal for fine-grained causal reasoning tasks.
In general, existing image datasets either lack explicit definitions of causal relations beyond visual representation or are too specific and simplified to enable comprehensive exploration of diverse causal relations. On the other hand, datasets in the causal research community, while rich in diverse causalities and clear causal definitions, lack corresponding visual representations, making them unsuitable for tasks involving causal reasoning in images. This disconnect makes it challenging to effectively evaluate AI systems’ abilities to reason about causality, thereby creating a significant bottleneck in advancing this field.
@@
  • Dataset. We introduce Causal3D, an innovative and comprehensive benchmark with 19 datasets that cover diverse causal structures, views, and backgrounds in realistic 3D scenes.
  • Evaluation. We implement a thorough evaluation of models on Causal3D, spanning from traditional causal algorithms to advanced LLMs and VLMs for images, offering a detailed analysis of the current state-of-the-art models on our benchmark.
  • Insights. We lay a strong foundation for advancing causal learning in CV by bridging the gap between these fields through our benchmark and provide key insights from our experimental observations.
@@ 2 Related Work
Causal Discovery from Tabular Data. Causal discovery is an important task in causal inference (Pearl, 2009), aiming to identify the causal relations from data. Multiple methods have been developed for this task and most of them focus on tabular data  (Wen et al., 2021; Tu et al., 2024; Wen et al., 2022; Cinquini et al., 2021; Russo and Toni, 2023). These methods mainly include constraint-based methods (e.g., PC (Spirtes et al., 2000b) and score-based methods (e.g., GES (Chickering, 2002)). Many of them are statistical methods (e.g., CAM (Bühlmann et al., 2014), LiNGAM (Yang et al., 2024)), which have good theoretical support but often suffer from strong assumptions, sensitivity, and scalability issues. Recently, deep learning-based methods have attracted lots of attention due to their improvement in these aspects. Among them, causal-TGAN (Wen et al., 2022), GraN-DAG (Lachapelle et al., 2020), DAG-GNN (Yu et al., 2019)) can model nonlinear causal relations in large datasets. Diffusion-based approaches (e.g., DiffAN (Sanchez et al., 2023)) improve robustness against noises while being computationally intensive and sensitive to hyperparameters.
@@
LLM-based Causal Discovery. Recent advances in LLMs have broadened their role in causal discovery (Ma, 2024; Jin et al., 2024; Liu et al., 2024b; Ban et al., 2023; Liu et al., 2024a; Wu et al., 2024; Wan et al., 2024; Shen et al., 2024). LLM-based causal discovery spans pairwise and full-graph discovery (Ma, 2024; Kıcıman et al., 2023; Jiralerspong et al., 2024), often leveraging prompts such as binary or multiple-choice selection and natural question-answering. Among them, many state-of-the-art models like ChatGPT 4o (Rawal et al., 2024) and Gemini-1.5 Pro (Carro et al., 2024) have been widely explored for causal inference. Besides, some causality-specific agents like Causal Copilot (Wang et al., 2024b) integrate LLMs for natural language-based causal queries. Despite its promising performance, LLMs still face key limitations, including difficulty in handling latent confounders and complex causal tasks.
@@
Causal Methods in CV. Causal inference in CV is essential for improving generalization and interpretability (Yang et al., 2021; Schölkopf, 2022). Many causal tasks have thus been widely explored in image data, one of them is causal representation learning, which aims to uncover disentangled and causally meaningful representations corresponding to high-level variables from data (Liu et al., 2022; Deng et al., 2022; Schölkopf et al., 2021). Many representative approaches in this area are based on generative models (e.g., CausalVAE (Yang et al., 2023), DEAR (Shen et al., 2022)). However, they often depend on strong assumptions (e.g., available annotation of high-level concepts) that may not always hold. Recently, explorations of VLMs in causal tasks under more general scenes have also attracted increasing attention (Wang et al., 2024a; Zhao et al., 2024).
@@
Causal Datasets in CV. Tabular data has long dominated causal inference research  (Runge et al., 2019a; Runge, 2018; Runge et al., 2019b; Spirtes et al., 2000a; Zheng et al., 2018). With the recent increasing need for causal studies in different data types and modalities, the community in CV and VLM has also placed more emphasis on causality. Recent years have witnessed the emergence of image and video datasets for causal reasoning (Zimmermann et al., 2021; Ates et al., 2022; Mao et al., 2022; Tung et al., 2024), bridging the gap between CV learning and causal reasoning. Despite these datasets addressing the absence of visual data encoded with causality, most of them still remain limited. They either focus on specific scenes, restricting the diversity of causal relations or lack rigorous causal definitions—such as explicit causal graphs and structured tabular records—to capture interactions among in-image variables. Consequently, a gap remains between traditional causal research and the study of causality in CV and VLM, as summarized in Tab. 1. This highlights the need for datasets that integrate explicit causal structures with both visual and tabular representations.
@@ 3 Causal3D: The Proposed Benchmark
We introduce Causal3D, a realistic 3D image dataset designed for casual learning. Aiming to bridge the gap between causal study and CV/VLM community, Causal3D is established with structured tabular data and tightly aligned visual representation. To build a comprehensive benchmark for evaluating models’ ability to uncover causality, Causal3D contains visual representations encoded with diverse causal relations among multiple variables. The dataset comprises two main components: Physically Consistent 3D Scenes, which simulate real-world settings to enhance the authenticity of the dataset, and Hypothetical 3D Scenes, introduced to diversify the causal relations represented in the dataset. To simulate realistic images in 3D scenes, we select or design causal relations to generate structured datasets and use Blender [^1] https://www.blender.org/ to render high-quality images, as shown in Fig. 2. Furthermore, by introducing various views and backgrounds, Causal3D presents the same scene in different surroundings, enhancing dataset diversity and contextual richness.
@@ 3.1 Dataset Components
Physically Consistent Scenes: In the physically consistent setting, Causal3D incorporates fundamental physical principles, such as light dynamics, magnetic fields, water pressure, and mechanics, to ensure realistic causal relations. To systematically vary the complexity of causal discovery, we curated 8 distinct scenes, each featuring 2 to 5 variables with unique causal structures. Each scene contains 10K samples. We provide visual overviews of each scene in Appendix A.
Hypothetical Scenes: Since causal relations in reality are often complex and difficult to observe, designing scenes with diverse causal graphs is challenging. In Causal3D, we also introduce additional hypothetical scenes that broaden the range of causal relations in our benchmark. Specifically, we explore causal relations under artificially defined hypothetical rules and synthesize causal relations among three fundamental 3D objects in Blender: sphere, cuboid, and cone. By defining specific dependencies among their variables (e.g., sphere radius, cuboid height), we construct both linear and non-linear causal relations across various graph configurations. This process yields 11 hypothetical scenes, each containing 10K samples. More details can be found in Appendix A.
@@ 3.3 Task Designs
Based on our dataset, we focus on three key causal tasks to evaluate state-of-the-art algorithms and models in discovering and leveraging causal relations across diverse causal structures and scenes. These tasks include:
Causal discovery from tabular data: This task focuses on identifying latent causal relations among variables using only tabular data. In this setting, we have high-level causal variable values recorded in the tabular data and do not rely on image information. It is evaluated based on the correctness of inferred causal structures across various datasets and underlying causal mechanisms.
Causal representation learning from images: This task aims to learn disentangled and causally meaningful representations from image data, meanwhile enabling models to capture underlying causal relations between high-level concepts. We use image data along with any additional information required by the models as input. Evaluation is based on the generated images after intervening on learned representations, assessing whether they accurately reflect the corresponding causal variables and causal relations.
Causal discovery & intervention from few images: This task focuses on uncovering causal relations and assessing intervention results with a limited number of images. Unlike traditional causal discovery methods that rely on large datasets, this task evaluates the ability of models to infer causal structures from a small set of images. Furthermore, we conduct causal interventions by manipulating a certain variable to observe its effect on the whole image. Causal intervention is often implemented with the do-operator $do(\cdot)$ (Pearl, 2009). Here, $do(X=x)$ means modifying a variable $X$ to a specific value $x$ while keeping all other influences unchanged. Intervention evaluation is based on whether the intervened images still remain consistent with the underlying causal relations.
@@ 4.1 Experiment Settings
Evaluation Metrics. We mainly use two metrics to quantify the experimental results: 1) F1 Score: F1 score is used to evaluate causal discovery results, which represents the harmonic mean of precision and recall of discovered causal relations. 2) Accuracy: Accuracy is used in causal intervention, which measures the fraction of consistent causal relations in the intervened image. Each experiment is repeated 10 times per setting, and the final results are obtained by averaging these results for robust evaluation.
Figure 3: Results of different causal discovery methods from tabular data on realistic scenes and hypothetical scenes.
@@ 4.2 Causal Discovery from Tabular Data
Evaluated Methods. We evaluate the performance of various causal discovery methods on tabular data, covering both traditional algorithms and emerging LLM-powered approaches. We assess 7 traditional methods, including CAM, NoTears, DAG GNN, DiffAN, PC, SCORE (Rolland et al., 2022), and GraN DAG. Additionally, we include the LLM-based method, Causal Copilot (Wang et al., 2024b).
@@ 4.3 Causal Representation Learning
Figure 4: Examples of the 4 causal representation learning model results. For each scene, the 3 columns represent: (1) original images, (2) Do(Cau): results after intervening on a “cause” variable, and (3) Do(Res): results after intervening on a “result” variable.
@@
Evaluation Strategy. Since many of these models require a (partial) causal graph as input or supervision, it would be unfair to directly evaluate them with regular causal discovery metrics we used in the previous subsection. Following the evaluation strategy of previous works (Yang et al., 2020; Shen et al., 2022; Komanduri et al., 2023; An et al., 2023), we apply interventions on each variable separately by modifying its corresponding causal representations, and then decode the modified latent representation back to generate an intervened image. By assessing whether the causally related variables in the image change accordingly, we can evaluate whether the model has effectively learned the causal relations within our dataset. As these models only accept 2D images, we render images of the 3D scene from a fixed viewpoint to serve as input. For CausalVAE, DEAR, and ICM-VAE, we use spring and seesaw for evaluation. However, for CDG-VAE, which requires a structured mask to distinguish each object (i.e., variable), meaning that each object must move within a specific region, spring and seesaw do not satisfy this requirement. Therefore, we select reflection and pendulum for evaluation.
@@
Results and Analysis. Intervention results are showcased in Fig. 4. For each given original image, we select a pair of “cause” and “result” variables to intervene with do-operation. For spring, the cause is the weight of the wooden block, and the result is the spring’s deformation. We first apply an intervention on the weight of the wooden block (cause), with the intervened image shown in the left-middle column. We observe that decreasing the block’s weight leads to a corresponding decrease in the spring’s deformation. This aligns with the causal relation, where the block’s weight directly influences the spring’s deformation. However, when we directly modify the spring’s deformation (result), as shown in the left-last column, the block’s weight remains unchanged. This confirms the unidirectional nature of causal relations, where the cause affects the effect, but not vice versa. The results for the seesaw scene exhibit a similar pattern. When we apply an intervention on the torque on the left side (cause), as shown in the right-middle column, the seesaw’s direction (result) reverses accordingly. However, when we directly intervene on the seesaw’s direction, the torque on both the left and right sides remains unchanged. Similar observations hold for reflection, where the incident light serves as the cause and the reflected light as the result, and for pendulum, where the pendulum’s angle acts as the cause, influencing the position and length of its shadow. Although these models may not perform optimally in certain scenes, exhibiting distortions in reconstructed images and incomplete disentanglement of attributes, they still capture some underlying causal relations. This indicates the rationality of our proposed benchmark and highlights the need for further research in achieving more effective causal representation learning from images.

### v2

@@ Abstract
True intelligence hinges on the ability to uncover and leverage hidden causal relations. Despite significant progress in AI and computer vision (CV), there remains a lack of benchmarks for assessing models’ abilities to infer latent causality from complex visual data. In this paper, we introduce Causal3D, a novel and comprehensive benchmark that integrates structured data (tables) with corresponding visual representations (images) to evaluate causal reasoning. Designed within a systematic framework, Causal3D comprises 19 3D-scene datasets capturing diverse causal relations, views, and backgrounds, enabling evaluations across scenes of varying complexity. We assess multiple state-of-the-art methods, including classical causal discovery, causal representation learning, and large/vision-language models (LLMs/VLMs). Our experiments show that as causal structures grow more complex without prior knowledge, performance declines significantly, highlighting the challenges even advanced methods face in complex causal scenarios. Causal3D serves as a vital resource for advancing causal reasoning in CV and fostering trustworthy AI in critical domains. Datasets link https://huggingface.co/datasets/LLDDSS/Causal3D_Dataset
@@ 1 Introduction
Figure 1: The proposed Causal3D dataset. We display 8 real-world scenes (11 hypothetical scenes are in the Appendix 6). We focus on 3 scenes: springs, parabolas, and water flow. 1) The blue block represents multi-view images of each scene, offering four different views. The first row shows virtual backgrounds, while the second row shows real backgrounds, with the same view in each column; 2) The green block provides textual descriptions; 3) The yellow block represents the causal graphs for each scene, along with the meanings of each variable in the graphs; 4) The pink block shows the structural equations (i.e., functions describing causal relations) for each scene. The bottom row briefly presents an overview of the remaining 5 real-world scenes and their corresponding causal graphs, including reflection, seesaw, convex lens, magnet, and pendulum. Detailed information on these 5 scenes can be found in the Appendix 6.
@@
Some existing datasets have been involved in causal studies in visual AI systems. However, these datasets often have significant limitations. For instance, oversimplified 2D hypothetical datasets (Yang et al., 2020) fail to capture the richness and complexity of real-world environments. Similarly, domain-specific datasets like the CelebA face dataset (Liu et al., 2015) are not designed for causal reasoning and lack the structural diversity required for comprehensive explorations. Recent datasets developed for vision- and multimodal-language models (VLMs/MLMs) (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Mao et al., 2022; Tung et al., 2024) have improved in complexity and realism but remain limited in explicitly representing causality and in offering diverse causal relations. The absence of clearly defined causal relations within visual representations, along with the lack of tabular records that are tightly aligned with these representations to provide guidance, makes such physics-aware VLMs/MLMs datasets suboptimal for fine-grained causal reasoning tasks. Especially, as interest in 3D data grows, causal learning in 3D settings introduces new challenges, opportunities, and insights. The complexity of realistic 3D scenes—encompassing lighting, texture, background, and viewpoint—can introduce spurious correlations and backdoor paths, making causal inference more difficult. At the same time, 3D environments provide multi-view consistency, allowing models to observe the same underlying causal relationships from diverse perspectives, which helps disentangle true causality from viewpoint-dependent features. This makes 3D datasets uniquely valuable for developing and evaluating robust, causally grounded models in realistic settings. However, this area remains underexplored.
In general, existing visual datasets either lack explicit definitions of causal relations beyond visual representation or are too specific and simplified to enable comprehensive exploration of diverse causal relations. On the other hand, datasets in the causal research community, while rich in diverse causalities and clear causal definitions, lack corresponding visual representations, making them unsuitable for tasks involving causal reasoning in images, not to mention complicated 3D scenarios. This disconnect makes it challenging to effectively advance and evaluate AI systems’ abilities of reliable reasoning, thereby creating a significant bottleneck in advancing this field.
@@
  • Dataset We introduce Causal3D, a novel and comprehensive benchmark consisting of 19 datasets that span a wide range of causal structures, viewpoints, and background variations within realistic 3D scenes. The full dataset will be released to support further research.
  • Evaluation We implement a thorough evaluation of models on Causal3D, spanning from traditional causal algorithms to advanced LLMs and VLMs for images, offering a detailed analysis of the current state-of-the-art models on our benchmark.
  • Insights We lay a strong foundation for advancing causal learning in CV by bridging the gap between these fields through our benchmark and provide key insights from our experimental observations.
@@ 2 Related work
Causal Discovery from Tabular Data Causal discovery is an important task in causal inference (Pearl, 2009), aiming to identify the causal relations from data. Multiple methods have been developed for this task and most of them focus on tabular data  (Wen et al., 2021; Tu et al., 2024; Wen et al., 2022; Cinquini et al., 2021; Russo and Toni, 2023). These methods mainly include constraint-based methods (e.g., PC (Spirtes et al., 2000b)) and score-based methods (e.g., GES (Chickering, 2002)). Many of them are statistical methods (e.g., CAM (Bühlmann et al., 2014), LiNGAM (Yang et al., 2024)), which have good theoretical support but often suffer from strong assumptions, sensitivity, and scalability issues.
Recently, deep learning-based methods have attracted lots of attention due to their improvement in these aspects. Among them, causal-TGAN (Wen et al., 2022), GraN-DAG (Lachapelle et al., 2020), DAG-GNN (Yu et al., 2019)) can model nonlinear causal relations in large datasets. Diffusion-based approaches (e.g., DiffAN (Sanchez et al., 2023)) improve robustness against noises while being computationally intensive and sensitive to hyperparameters.
@@
LLM-based Causal Discovery Recent advances in LLMs have broadened their role in causal discovery (Ma, 2024; Jin et al., 2024; Liu et al., 2024b; Ban et al., 2023; Liu et al., 2024a; Wu et al., 2024; Wan et al., 2024; Shen et al., 2024). LLM-based causal discovery spans pairwise and full-graph discovery (Ma, 2024; Kıcıman et al., 2023; Jiralerspong et al., 2024), often leveraging prompts such as binary or multiple-choice selection and natural question-answering. Among them, many state-of-the-art models like ChatGPT 4o (Rawal et al., 2024) and Gemini-1.5 Pro (Carro et al., 2024) have been widely explored for causal inference. Besides, some causality-specific agents like Causal Copilot (Wang et al., 2024b) integrate LLMs for natural language-based causal queries. Despite its promising performance, LLMs still face key limitations, including difficulty in handling latent confounders and complex causal tasks.
@@
Causal Methods in CV Causal inference in CV is essential for improving generalization and interpretability (Yang et al., 2021; Schölkopf, 2022). Many causal tasks have thus been widely explored in image data, one of them is causal representation learning, which aims to uncover disentangled and causally meaningful representations corresponding to high-level variables from data (Liu et al., 2022; Deng et al., 2022; Schölkopf et al., 2021). Many representative approaches in this area are based on generative models (e.g., CausalVAE (Yang et al., 2023), DEAR (Shen et al., 2022)). However, they often depend on strong assumptions (e.g., available annotation of high-level concepts) that may not always hold. Recently, explorations of VLMs in causal tasks under more general scenes have also attracted increasing attention (Wang et al., 2024a; Zhao et al., 2024).
@@
Causal Datasets in CV Tabular data has long dominated causal inference research  (Runge et al., 2019a; Runge, 2018; Runge et al., 2019b; Spirtes et al., 2000a; Zheng et al., 2018). With the recent increasing need for causal studies in different data types and modalities, the community in CV and VLM has also placed more emphasis on causality. Recent years have witnessed the emergence of image and video datasets for causal reasoning (Zimmermann et al., 2021; Ates et al., 2022; Mao et al., 2022; Tung et al., 2024), bridging the gap between CV learning and causal reasoning. Despite these datasets addressing the absence of visual data encoded with causality, most of them still remain limited. They either focus on specific scenes, restricting the diversity of causal relations or lack rigorous causal definitions—such as explicit causal graphs and structured tabular records—to capture interactions among in-image variables. Consequently, a gap remains between traditional causal research and the study of causality in CV and VLM, as summarized in Tab. 1. This highlights the need for datasets that integrate explicit causal structures with both visual and tabular representations.
@@ 3 Causal3D: the proposed benchmark
We introduce Causal3D, a realistic 3D image dataset designed for casual learning from observational visual data. Aiming to bridge the gap between causal study and CV/VLM community, Causal3D is established with structured tabular data and tightly aligned visual representation. To build a comprehensive benchmark for evaluating models’ ability to uncover causality, Causal3D contains visual representations encoded with diverse causal relations among multiple variables. The dataset comprises two main components: Physically Consistent 3D Scenes, which simulate real-world settings to enhance the authenticity of the dataset, and Hypothetical 3D Scenes, introduced to diversify the causal relations represented in the dataset. To simulate realistic images in 3D scenes, we select or design causal relations to generate structured datasets and use Blender [^1] https://www.blender.org/ to render high-quality images, as shown in Fig. 2. Furthermore, by introducing various views and backgrounds, Causal3D presents the same scene in different surroundings, enhancing dataset diversity and contextual richness.
@@ 3.1 Dataset components
Physically Consistent Scenes
In the physically consistent setting, Causal3D incorporates fundamental physical principles, such as light dynamics, magnetic fields, water pressure, and mechanics, to ensure realistic causal relations. To systematically vary the complexity of causal discovery, we curated 8 distinct scenes, each featuring 2 to 5 variables with unique causal structures. Each scene contains 10K samples. We provide visual overviews of each scene in Appendix 6.
Hypothetical Scenes
Since causal relations in reality are often complex and difficult to observe, designing scenes with diverse causal graphs is challenging. In Causal3D, we also introduce additional hypothetical scenes that broaden the range of causal relations in our benchmark. Specifically, we explore causal relations under artificially defined hypothetical rules and synthesize causal relations among three fundamental 3D objects in Blender: sphere, cuboid, and cone. By defining specific dependencies among their variables (e.g., sphere radius, cuboid height), we construct both linear and non-linear causal relations across various graph configurations. This process yields 11 hypothetical scenes, each containing 10K samples. More details can be found in Appendix 6.
@@ 3.3 Task designs
Based on our dataset, we focus on three key causal tasks to evaluate state-of-the-art algorithms and models in discovering and leveraging causal relations across diverse causal structures and scenes. We focus on the widely adopted setting of causal learning from observational data. These tasks include:
Causal discovery from tabular data
This task focuses on identifying latent causal relations among variables using only tabular data. In this setting, we have high-level causal variable values recorded in the tabular data and do not rely on image information. It is evaluated based on the correctness of inferred causal structures across various datasets and underlying causal mechanisms.
Causal representation learning from images
This task aims to learn disentangled and causally meaningful representations from images, meanwhile enabling models to capture underlying causal relations between high-level concepts. We use image data along with any additional information required by the models as input. Evaluation is based on generated images after intervening on learned representations, assessing whether they accurately reflect the corresponding causal variables and causal relations.
Causal discovery & intervention from few images
This task focuses on uncovering causal relations and assessing intervention results with a limited number of images. Unlike traditional causal discovery methods that rely on large datasets, this task evaluates the ability of models to infer causal structures from a small set of images. Furthermore, we conduct causal interventions by manipulating a certain variable to observe its effect on the whole image. Causal intervention is often implemented with the do-operator $do(\cdot)$ (Pearl, 2009). Here, $do(X=x)$ means modifying a variable $X$ to a specific value $x$ while keeping all other influences unchanged. Intervention evaluation is based on whether the intervened images still remain consistent with the underlying causal relations.
@@ 4.1 Experiment settings
Evaluation Metrics. We mainly use two metrics to quantify the experimental results: 1) F1 Score is used to evaluate causal discovery results, which represents the harmonic mean of precision and recall of discovered causal relations. 2) Accuracy is used in causal intervention, which measures the fraction of consistent causal relations in the intervened image. Each experiment is repeated 10 times per setting, and the final results are obtained by averaging these results for robust evaluation.
Figure 3: Results of different causal discovery methods from tabular data on realistic/hypothetical scenes.
@@ 4.2 Causal discovery from tabular data
Evaluated Methods. We evaluate the performance of various causal discovery methods on tabular data, covering both traditional algorithms and emerging LLM-powered approaches. We assess 7 traditional methods: CAM, NoTears, DAG GNN, DiffAN, PC, SCORE (Rolland et al., 2022), and GraN DAG; 1 LLM-based method: Causal Copilot (Wang et al., 2024b). All experiments were conducted on an NVIDIA RTX 4090 GPU.
@@ 4.3 Causal representation learning
Figure 4: Examples of the 4 causal representation learning model results. In each scene, the 3 columns show: 1) original images, 2) Do(Cau): results after intervening on a “cause” variable, and 3) Do(Res): after intervening on a “result” variable.
@@
Evaluation Strategy. Since many of these models require a (partial) causal graph as input or supervision, it would be unfair to directly evaluate them with regular causal discovery metrics we used in the previous subsection. Following the evaluation strategy of previous works (Yang et al., 2020; Shen et al., 2022; Komanduri et al., 2023; An et al., 2023), we apply interventions on each variable separately by modifying its corresponding causal representations, and then decode the modified latent representation back to generate an intervened image. By assessing whether the causally related variables in the image change accordingly, we can evaluate whether the model has effectively learned the causal relations within our dataset.
Since these models only accept 2D images, we render the 3D scenes from a fixed viewpoint as inputs. For CausalVAE, DEAR, and ICM-VAE, we use spring and seesaw for evaluation. However, for CDG-VAE, which requires a structured mask to distinguish each object (i.e., variable), meaning that each object must move within a specific region, spring and seesaw do not satisfy this requirement. Therefore, we select reflection and pendulum for evaluation.
@@
Results and Analysis. Intervention results are showcased in Fig. 4. For each given original image, we select a pair of “cause” and “result” variables to intervene with do-operation. For spring, the cause is the weight of the wooden block, and the result is the spring’s deformation. We first apply an intervention on the weight of the wooden block (cause), with the intervened image shown in the left-middle column. We observe that decreasing the block’s weight leads to a corresponding decrease in the spring’s deformation. This aligns with the causal relation, where the block’s weight directly influences the spring’s deformation. However, when we directly modify the spring’s deformation (result), as shown in the left-last column, the block’s weight remains unchanged. This confirms the unidirectional nature of causal relations, where the cause affects the effect, but not vice versa. The results for the seesaw scene exhibit a similar pattern. When we apply an intervention on the torque on the left side (cause), as shown in the right-middle column, the seesaw’s direction (result) reverses accordingly. However, when we directly intervene on the seesaw’s direction, the torque on both the left and right sides remains unchanged. Similar observations hold for reflection, where the incident light serves as the cause and the reflected light as the result, and for pendulum, where the pendulum’s angle acts as the cause, influencing the position and length of its shadow.
Although these models may not perform optimally in certain scenes, exhibiting distortions in reconstructed images and incomplete disentanglement of attributes, they still capture some underlying causal relations. This indicates the rationality of our proposed benchmark and highlights the need for further research in achieving more effective causal representation learning from images.